# Strong time-periodic solutions to a multilayered fluid-structure interaction problem

### v1

@@ Abstract.
This article studies a nonlinearly coupled, time-periodic fluid-structure interaction (FSI) problem involving the Navier-Stokes equations for an incompressible, viscous fluid coupled with a multilayered elastic structure consisting of a thin damped shell and a thick viscoelastic component. The main objective is to establish the existence of a strong time-periodic solution. To address the full nonlinear problem, we employ a fixed point argument, which relies on a detailed analysis of the linearized system together with appropriate nonlinear estimates. The linearized problem is analyzed by employing the Arendt-Bu theorem on maximal periodic $\mathrm{L}^{p}$-regularity, which requires several new analytical ingredients including a careful treatment of the thick structural layer, a refined lifting procedure, a decoupling strategy to establish $\mathcal{R}$-sectoriality of the associated FSI operator and a careful spectral analysis. To the best of our knowledge, this is the first result proving the existence of strong time-periodic solutions for a multilayered fluid-structure interaction problem.
@@ 1. Introduction
The present paper establishes for the first time the existence of a strong time-periodic solution to a multilayered FSI problem subject to time-periodic forcing terms. In fact, we consider the nonlinearly coupled interaction problem of a viscous, incompressible Newtonian fluid with a multilayered structure consisting of a thin plate-type and a thick (visco-)elastic layer. Here the term nonlinear coupling means that the fluid-structure interface is not fixed, but it is evolving in time, giving rise to a moving boundary problem. The model will be introduced in detail in Section 2 and can be found in (2.17). The motivation for studying multilayered FSI problems is strongly rooted in biomedical applications, where many physiological structures, such as blood vessels, are composed of multiple elastic and viscoelastic layers. We refer here for example to plaque growth in arteries, which is a key factor in atherosclerosis and whose modeling and numerical simulations can e. g. be found in [24, 52].
@@
Concerning strong solutions to FSI problems with elastic structures located at the fluid boundary, a first local-in-time well-posedness result for fluid-structure systems involving viscoelastic beams and small initial data was established in the foundational work [3]. Remarkably, Grandmont and Hillairet [26] proved the global-in-time strong well-posedness of such FSI problems in a 2D fluid-1D beam setting for large data. Strong solutions to FSI problems involving nonlinear elastic Koiter shells were studied in [14, 39], while approaches to FSI problems of incompressible fluids and damped (thin) structures via $\mathrm{L}^{p}$-theory have been carried out by Denk and Saal [18] and Maity and Takahashi [41]. Let us emphasize that none of these works addresses the situation of a multilayered structure.
More recently, attention has also turned to FSI problems of greater complexity, including the case of compressible or heat-conducting fluids coupled with beams and shells, see e. g. [8, 40, 38]. In [33], Kuan, Čanić and Muha studied a multilayered FSI problem, where the thick layer is assumed to be poroelastic.
The study of time-periodic problems is a classical topic in mathematical fluid mechanics. We refer here for instance to the pioneering articles by Serrin [55], Judovič [27] and Prodi [53] as well as the works of Kozono and Nakao [31], Galdi and Sohr [22] or Geissert, Hieber and Nguyen [23]. On the other hand, time-periodicity plays a central role in modeling phenomena where cyclic behavior is inherent, especially in biomedical applications. For example, the rhythmic nature of blood flow, governed by the heart’s repetitive pumping, naturally leads to models with periodic forcing. However, despite their relevance, FSI problems in the time-periodic setting remain underexplored. Several recent studies have taken important steps toward overcoming the associated analytical difficulties. In the case of incompressible fluids interacting with (single) elastic structures, results on strong solutions have been obtained by Casanova [11] as well as Celik and Kyed [12]. The existence of weak solutions to FSI problems in time-periodic settings has been rigorously investigated in [44, 45] and [43], where the latter work addresses a multilayered FSI problem. For time-periodic solutions to FSI problems involving different fluid models, such as compressible ones, and structural configurations, we also refer to [28, 51, 32]. Time-periodic weak solutions to parabolic-hyperbolic systems have also been investigated in [46], while time-periodic fluid-rigid body interaction problems have been explored e. g. by Bonheure and Galdi [6].
@@
Let us elaborate on the main contributions and methodological advances of this article. This work appears to be the first to establish the existence of strong solutions for a multilayered FSI problem. The presence of the thick structural layer introduces substantial analytical challenges, requiring significant extensions beyond existing FSI theory. A key technical contribution is the bounded $\mathcal{H}^{\infty}$-calculus of the operator matrix associated with the damped wave system (see Subsection 4.2), which, to the best of our knowledge, is new and may be of independent interest. The presence of the thick layer also necessitates a further step in the lifting procedure, and the resulting multilayered fluid-structure operator is considerably more complex than in the case of a single thin layer as in [41]. In particular, to tackle the strongly coupled nature of the present multilayered problem, a more intricate decoupling approach tailored to the functional analytic setting of the damped wave problem is required. As far as we are aware, this is the first time such a technique has been employed in the context of boundary-coupled FSI problems with elastic structures, though it may prove useful in a broader class of FSI models. Furthermore, the damping in the thick structural layer renders the spectral analysis substantially more delicate. In particular, the determination of the essential spectrum requires a careful decomposition of the decoupled operator matrix.
@@
This article has the following outline. In Section 2, we introduce the time-periodic multilayered FSI problem and present the main result on the existence of a strong time-periodic solution for small periodic forces. Section 3 is dedicated to the transformation to a fixed fluid domain. The central Section 4 discusses the linear theory for the time-periodic problem, including the damped wave problem, the reformulation in operator form, and the maximal $\mathrm{L}^{p}$-regularity and invertibility of the resulting operator matrix. In Section 5, we establish the nonlinear estimates and complete the proof of the main result.
2. The multilayered FSI problem and main result
In this section, we first introduce the multilayered FSI problem and then state the main result on the existence of a strong time-periodic solution to this problem under the action of small time-periodic forces.
We consider the 3D viscous, incompressible Navier-Stokes equations, for the fluid. Due to the nonlinear coupling, the interface separating the fluid and the thick structural layer is generally not fixed, but it can move when time changes. As a result, the fluid domain $\Omega_{\mathrm{f}}(t)$ depends on the plate displacement, i. e., it is of the form $\Omega_{\mathrm{f}}(\eta(t))$, for $\eta$ representing the displacement of the plate. For simplicity, we will usually write $\Omega_{\mathrm{f}}(t)$ in the sequel. As we investigate the time-periodic problem, for a fixed time period $T>0$, we study the multilayered FSI problem on a time interval of the form $(0,T)$. For the fluid velocity $u^{\mathcal{F}}$ and pressure $\pi^{\mathcal{F}}$ on the moving domain, the resulting fluid subproblem reads as
(2.1) $\partial_{t}u^{\mathcal{F}}+(u^{\mathcal{F}}\cdot\nabla)u^{\mathcal{F}}=\operatorname{div}\sigma_{\mathrm{f}}(u^{\mathcal{F}},\pi^{\mathcal{F}})\enspace\text{and}\enspace\operatorname{div}u^{\mathcal{F}}=0,\enspace\text{for}\enspace t\in(0,T),\enspace x\in\Omega_{\mathrm{f}}(t),$
where
(2.2) $\sigma_{\mathrm{f}}(u^{\mathcal{F}},\pi^{\mathcal{F}})=2\mu_{\mathrm{f}}\mathbb{D}(u^{\mathcal{F}})-\pi^{\mathcal{F}}\operatorname{Id}_{3},\enspace\text{with}\enspace\mathbb{D}(u^{\mathcal{F}})\coloneqq\frac{1}{2}\bigl{(}\nabla u^{\mathcal{F}}+(\nabla u^{\mathcal{F}})^{\top}\bigr{)},$
Concerning the geometry of the problem, we suppose that the boundary of the (moving) fluid domain consists of a moving part $\Gamma_{\eta}(t)$, representing the interface with the thin structure, as well as a rigid part $\Gamma_{0}$. More precisely, we assume that the domain of the thin structure $\omega\subset\mathbb{R}^{2}$ is a smooth subset of $\mathbb{R}^{2}$, and this in turn implies that $\Gamma_{\eta}(t)=\{(s,\eta(s)):s\in\omega\}$. It follows that $\Gamma_{0}=\partial\Omega_{\mathrm{f}}(t)\setminus\Gamma_{\eta}(t)$.
For the plate, we assume that its motion is governed by a damped plate equation in $\omega$. After rescaling by the (thin) structure thickness and density, and for $f_{\mathrm{p}}$ denoting the force density acting on the structure and to be specified in the sequel, the linearly (visco-)elastic Koiter shell model is given by
$\partial_{tt}\eta+C_{0}\eta-C_{1}\Delta_{s}\eta+C_{2}\Delta_{s}^{2}\eta+D_{0}\partial_{t}\eta-D_{1}\partial_{t}\Delta_{s}\eta+D_{2}\partial_{t}\Delta_{s}^{2}\eta=f_{\mathrm{p}},\enspace\text{for}\enspace t\in(0,T),\enspace s\in\omega.$
Here $C_{i}\geq 0$ and $D_{i}\geq 0$ are the (rescaled) material constants corresponding to structural elasticity and viscosity, respectively, and the subscript s indicates that the respective differential operators act with respect to $\omega$, i. e., these are horizontal operators. In the sequel, we will take into account the situation of $C_{2}=D_{1}=1$ and $C_{0}=C_{1}=D_{0}=D_{2}=0$, i. e.,
(2.3) $\partial_{tt}\eta+\Delta_{s}^{2}\eta-\partial_{t}\Delta_{s}\eta=f_{\mathrm{p}},\enspace\text{for}\enspace t\in(0,T),\enspace s\in\omega.$
Moreover, we suppose that the deformed plate does not have contact with the other part of the fluid domain, so $\Gamma_{0}\cap\Gamma_{\eta}(t)=\emptyset$ for all $t\in(0,T)$.
In the following, in order to ensure strong ellipticity of the associated Lamé operator, we assume that
(2.5) $\mu_{\mathrm{s}}>0\enspace\text{and}\enspace\mu_{\mathrm{s}}+\lambda_{\mathrm{s}}>0.$
The structural problem is already defined in the Lagrangian framework, and we denote the associated domain by $\Omega_{\mathrm{s}}$. By the Lagrangian formulation, this domain does not depend on time or the displacement of the plate. The boundary of $\Omega_{\mathrm{s}}$ consists of two parts, namely the flat interface $\Gamma_{\mathrm{s}}$ as introduced above and the remaining part $\partial\Omega_{\mathrm{s}}\setminus\Gamma_{\mathrm{s}}$. The equations of linear (visco-)elasticity read as
(2.6) $\partial_{tt}d=\operatorname{div}\sigma_{\mathrm{s}}(d,\partial_{t}d),\enspace\text{for}\enspace t\in(0,T),\enspace x\in\Omega_{\mathrm{s}}.$
(2.7) $u^{\mathcal{F}}(t,s,\eta(t,s))=\partial_{t}\eta(t,s)\mathrm{e}_{3},\enspace\text{for}\enspace t\in(0,T),\enspace s\in\omega.$
For $x=(s,z)$, the continuity of displacement between the thick and the thin structure translates as
(2.8) $d(t,x)=\eta(t,s)\mathrm{e}_{3},\enspace\text{for}\enspace t\in(0,T),\enspace x\in\Gamma_{\mathrm{s}}.$
The dynamic coupling condition describes the balance of forces at the fluid-structure interface $\Gamma_{\eta}(t)$, and it enters in the shape of the forcing term $f_{\mathrm{p}}$ in (2.3), with forces on the plate exerted by the fluid and the thick structure. With the outer unit normal vector
$n_{\Gamma_{\eta}(t)}=\frac{1}{\sqrt{1+|\nabla_{s}\eta|^{2}}}\binom{-\nabla_{s}\eta}{1}$
to $\Gamma_{\eta}(t)$, and with the Cauchy stress tensor $\sigma_{\mathrm{f}}$ from (2.2) and the Piola-Kirchhoff stress tensor $\sigma_{\mathrm{s}}$ from (2.4), the dynamic coupling conditions can be rephrased as
(2.9) $f_{\mathrm{p}}=-\sqrt{1+|\nabla_{s}\eta|^{2}}\left.\left(\sigma_{\mathrm{f}}(u^{\mathcal{F}},\pi^{\mathcal{F}})n_{\Gamma_{\eta}(t)}\right)\right|_{\Gamma_{\eta}(t)}\cdot\mathrm{e}_{3}+\left.(\sigma_{\mathrm{s}}(d,\partial_{t}d)\mathrm{e}_{3})\right|_{\Gamma_{\mathrm{s}}}\cdot\mathrm{e}_{3}.$
We now prescribe the boundary conditions for the fluid velocity $u^{\mathcal{F}}$ at the rigid part $\Gamma_{0}$ of the fluid boundary, for the plate at the boundary $\partial\omega$ of its domain $\omega$ and for the displacement $d$ of the thick structure at the non-interface part $\partial\Omega_{\mathrm{s}}\setminus\Gamma_{\mathrm{s}}$ of the boundary of the thick layer. Here $u^{\mathcal{F}}$ and $d$ satisfy homogeneous Dirichlet boundary conditions, while the plate is clamped, i. e.,
(2.10) $u^{\mathcal{F}}(t,x)$ $=0,\enspace\text{for}\enspace t\in(0,T),\enspace x\in\Gamma_{0},\enspace\eta=\nabla_{s}\eta\cdot n_{\omega}=0,\enspace\text{for}\enspace t\in(0,T),\enspace s\in\partial\omega,\enspace\text{and}\enspace$
$d(t,x)$ $=0,\enspace\text{for}\enspace t\in(0,T),\enspace x\in\partial\Omega_{\mathrm{s}}\setminus\Gamma_{\mathrm{s}},$
where $n_{\omega}$ represents the unit exterior normal vector to $\partial\omega$. Finally, the time-periodicity condition is
(2.11) $u^{\mathcal{F}}(0,\cdot)$ $=u^{\mathcal{F}}(T,\cdot),\enspace\text{in}\enspace\Omega_{\mathrm{f}}(\eta(0)),\enspace\eta(0,\cdot)=\eta(T,\cdot)\enspace\text{and}\enspace\partial_{t}\eta(0,\cdot)=\partial_{t}\eta(T,\cdot),\enspace\text{in}\enspace\omega,$
$d(0,\cdot)$ $=d(T,\cdot)\enspace\text{and}\enspace\partial_{t}d(0,\cdot)=\partial_{t}d(T,\cdot),\enspace\text{in}\enspace\Omega_{\mathrm{s}}.$
Moreover, we consider suitable forcing terms $f\colon(0,T)\times\Omega_{\mathrm{f}}\to\mathbb{R}^{3}$ in the fluid equation, $g\colon(0,T)\times\omega\to\mathbb{R}$ in the plate equation and $h\colon(0,T)\times\Omega_{\mathrm{s}}\to\mathbb{R}^{3}$ in the equations of linear (visco-)elasticity.
For the sake of a more compact representation, we summarize the interaction problem consisting of the respective fluid, plate and linearly (visco-)elastic subproblem (2.1), (2.3) and (2.6) as well as the coupling conditions (2.7)–(2.9), the boundary conditions (2.10) and the time-periodicity conditions (2.11). At this stage, let us also remark that we take into account the volume constraint
(2.12) $\int_{\omega}\eta\,\mathrm{d}s=0$
in the following. Thus, we introduce the space $\mathrm{L}_{\mathrm{m}}^{q}(\omega)$ of functions with average zero in $\mathrm{L}^{q}(\omega)$, so
(2.13) $\mathrm{L}_{\mathrm{m}}^{q}(\omega)\coloneqq\left\{f\in\mathrm{L}^{q}(\omega):\int_{\omega}f\,\mathrm{d}s=0\right\}.$
We define $\mathrm{L}_{\mathrm{m}}^{q}(\Omega_{\mathrm{f}})$ likewise. In addition, we invoke the orthogonal projection $P_{\mathrm{m}}\colon\mathrm{L}^{q}(\omega)\to\mathrm{L}_{\mathrm{m}}^{q}(\omega)$ onto the space of functions with average zero. It is given by
(2.14) $P_{\mathrm{m}}f\coloneqq f-\frac{1}{|\omega|}\int_{\omega}f\,\mathrm{d}s,\enspace\text{for}\enspace f\in\mathrm{L}^{q}(\omega).$
$\quad+P_{\mathrm{m}}\left(\left.(\sigma_{\mathrm{s}}(d,\partial_{t}d)\mathrm{e}_{3})\right|_{\Gamma_{\mathrm{s}}}\cdot\mathrm{e}_{3}\right)$
and
$\quad-\int_{\omega}\left.(\sigma_{\mathrm{s}}(d,\partial_{t}d)\mathrm{e}_{3})\right|_{\Gamma_{\mathrm{s}}}\cdot\mathrm{e}_{3}\,\mathrm{d}s.$
The resulting problem, with (2.15) and (2.12) replacing (2.3), for $f_{\eta}\coloneqq f\circ X_{\eta}$ for a diffeomorphism $X_{\eta}\colon(0,T)\times\Omega_{\mathrm{f}}\to(0,T)\times\Omega_{\mathrm{f}}(t)$ made precise in Section 3, and with $\mathcal{J}_{\eta}\coloneqq\sqrt{1+|\nabla_{s}\eta|^{2}}$, reads as
(2.17) $\partial_{t}u^{\mathcal{F}}+(u^{\mathcal{F}}\cdot\nabla)u^{\mathcal{F}}&=\operatorname{div}\sigma_{\mathrm{f}}(u^{\mathcal{F}},\pi^{\mathcal{F}})+f_{\eta},\enspace\operatorname{div}u^{\mathcal{F}}=0,&&t\in(0,T),\enspace x\in\Omega_{\mathrm{f}}(t),\\ \partial_{tt}\eta+P_{\mathrm{m}}\Delta_{s}^{2}\eta-\partial_{t}\Delta_{s}\eta&=-P_{\mathrm{m}}\left(\mathcal{J}_{\eta}\left.\left(\sigma_{\mathrm{f}}(u^{\mathcal{F}},\pi^{\mathcal{F}})n_{\Gamma_{\eta}(t)}\right)\right|_{\Gamma_{\eta}(t)}\cdot\mathrm{e}_{3}\right)\\ &\qquad+P_{\mathrm{m}}\left(\left.(\sigma_{\mathrm{s}}(d,\partial_{t}d)\mathrm{e}_{3})\right|_{\Gamma_{\mathrm{s}}}\cdot\mathrm{e}_{3}\right)+P_{\mathrm{m}}g,&&t\in(0,T),\enspace s\in\omega,\\ \int_{\omega}\eta(t,s)\,\mathrm{d}s&=0,&&t\in(0,T),\\ \partial_{tt}d&=\operatorname{div}\sigma_{\mathrm{s}}(d,\partial_{t}d)+h,&&t\in(0,T),\enspace x\in\Omega_{\mathrm{s}},\\ u^{\mathcal{F}}(t,s,\eta(t,s))&=\partial_{t}\eta(t,s)\mathrm{e}_{3},&&t\in(0,T),\enspace s\in\omega,\\ d(t,x)&=\eta(t,s)\mathrm{e}_{3},&&t\in(0,T),\enspace x\in\Gamma_{\mathrm{s}},\\ u^{\mathcal{F}}(t,x)&=0,&&t\in(0,T),\enspace x\in\Gamma_{0},\\ d(t,x)&=0,&&t\in(0,T),\enspace x\in\partial\Omega_{\mathrm{s}}\setminus\Gamma_{\mathrm{s}},\\ \eta=\nabla_{s}\eta\cdot n_{\omega}&=0,&&t\in(0,T),\enspace s\in\partial\omega,\\ u^{\mathcal{F}}(0,\cdot)&=u^{\mathcal{F}}(T,\cdot),&&\enspace\text{in}\enspace\Omega_{\mathrm{f}}(\eta(0)),\\ \eta(0,\cdot)&=\eta(T,\cdot)\enspace\text{and}\enspace\partial_{t}\eta(0,\cdot)=\partial_{t}\eta(T,\cdot),&&\enspace\text{in}\enspace\omega,\\ d(0,\cdot)&=d(T,\cdot)\enspace\text{and}\enspace\partial_{t}d(0,\cdot)=\partial_{t}d(T,\cdot),&&\enspace\text{in}\enspace\Omega_{\mathrm{s}}.\end{aligned}\right.$
Before stating the corresponding theorem, we require some preparation in terms of notation. More precisely, let $\eta\in\mathrm{L}^{p}(0,T;\mathrm{W}^{4,q}(\omega))\cap\mathrm{W}^{2,p}(0,T;\mathrm{L}^{q}(\omega))$ such that $\eta=0$ and $\nabla_{s}\eta\cdot n_{\omega}=0$ is valid on $\partial\omega$, and we have $\Gamma_{0}\cap\Gamma_{\eta(t)}=\emptyset$. In Section 3, the existence of a map $X=X_{\eta}$ being a $\mathrm{C}^{1}$-diffeomorphism from $\Omega_{\mathrm{f}}$ onto $\Omega_{\mathrm{f}}(t)$ and satisfying $X\in\mathrm{L}^{p}\bigl{(}0,T;\mathrm{W}^{2,q}(\Omega_{\mathrm{f}})^{3}\bigr{)}\cap\mathrm{W}^{2,p}\bigl{(}0,T;\mathrm{L}^{q}(\Omega_{\mathrm{f}})^{3}\bigr{)}$ is shown. For the above $T>0$ as well as $p$, $q\in(1,\infty)$, we then define the function spaces

### v2

# Strong time-periodic solutions for a multilayered fluid-structure interaction system with nonlinear coupling

@@ Abstract.
We investigate a time-periodic fully three-dimensional fluid-structure interaction system in which the Navier-Stokes equations for an incompressible viscous fluid are coupled with a multilayered elastic structure composed of a damped thin linear plate and a thick viscoelastic layer. The coupling is nonlinear, meaning that it is on a moving interface that is not known a priori, rendering the problem a moving-domain problem. We prove the existence of strong time-periodic solutions. The proof relies on a fixed point argument, combining sharp nonlinear estimates with a detailed analysis of the linearized system. The linearized problem is analyzed by employing the Arendt-Bu theorem on maximal periodic $\mathrm{L}^{p}$-regularity, which requires several new analytical ingredients including a refined lifting procedure, a decoupling strategy establishing $\mathcal{R}$-sectoriality of the coupled operator, a careful treatment of the thick structural layer, and a spectral analysis adapted to the multilayered setting. This provides the first strong time-periodic existence result for multilayered fluid-structure interaction systems, and the methods are expended to extend more broadly to nonlinear coupled PDEs on moving domains with periodic forcing.
@@ 1. Introduction
The present paper establishes for the first time the existence of a strong time-periodic solution to a fully three-dimensional multilayered FSI problem subject to time-periodic forcing terms. In fact, we consider the nonlinearly coupled interaction problem of a three-dimensional viscous, incompressible Newtonian fluid with a multilayered structure consisting of a two-dimensional thin plate-type and a three-dimensional thick (visco-)elastic layer. Here the term nonlinear coupling means that the fluid-structure interface is not fixed, but it is evolving in time, giving rise to a moving boundary problem. The model will be introduced in detail in Section 3 and can be found in (3.16). The motivation for studying multilayered FSI problems is strongly rooted in biomedical applications, where many physiological structures, such as blood vessels, are composed of multiple elastic and viscoelastic layers. We refer here for example to plaque growth in arteries, which is a key factor in atherosclerosis and whose modeling and numerical simulations can, e.g., be found in [31, 59]. Another application of multilayered FSI problems concerns bioartificial organ design, see for instance [66].
@@
Concerning strong solutions to FSI problems with elastic structures located at the fluid boundary, a first local-in-time well-posedness result for fluid-structure systems involving viscoelastic beams and small initial data was established in the foundational work [7]. Remarkably, Grandmont and Hillairet [33] proved the global-in-time strong well-posedness of such FSI problems in a 2D fluid-1D beam setting for large data. Strong solutions to FSI problems involving nonlinear elastic Koiter shells were studied in [20, 46], while approaches to FSI problems of incompressible fluids and damped (thin) structures via $\mathrm{L}^{p}$-theory have been carried out by Denk and Saal [24] and Maity and Takahashi [48]. Let us emphasize that all of these works concern the initial value problem, and none of these works addresses the situation of a multilayered structure.
There are contributions on multilayered and coupled PDE systems, including studies on multilayered Lamé-heat systems or heat-wave-wave interactions, which provide insight into stability, decay, and spectral properties of multilayered coupled systems in linear settings [6, 4, 3, 5]. Multilayered FSI problems involving poroelastic structures were studied in the linearly coupled case in [10] and in the nonlinearly coupled case in [40, 12]. Recently, attention has also turned to FSI problems of greater complexity, including the case of compressible or heat-conducting fluids coupled with beams and shells, see, e.g., [14, 47, 45].
The study of time-periodic problems is a classical topic in mathematical fluid mechanics. We refer here for instance to the pioneering articles by Serrin [62], Judovič [34] and Prodi [60] as well as the works of Kozono and Nakao [38], Galdi and Sohr [28] or Geissert, Hieber and Nguyen [30]. On the other hand, time-periodicity plays a central role in modeling phenomena where cyclic behavior is inherent, especially in biomedical applications. For example, the rhythmic nature of blood flow, governed by the heart’s repetitive pumping, naturally leads to models with periodic forcing. However, despite their relevance, FSI problems in the time-periodic setting remain underexplored. Several recent studies have taken important steps toward overcoming the associated analytical difficulties. In the case of incompressible fluids interacting with (single) elastic structures, results on strong solutions have been obtained by Casanova [17] as well as Celik and Kyed [18]. The existence of weak solutions to FSI problems in time-periodic settings has been rigorously investigated in [51, 52] and [50], where the latter work addresses a multilayered FSI problem. For time-periodic solutions to FSI problems involving different fluid models, such as compressible ones, and structural configurations, we also refer to [35, 58, 39]. Time-periodic weak solutions to parabolic-hyperbolic systems have also been investigated in [53], while time-periodic fluid-rigid body interaction problems have been explored, e.g., by Bonheure and Galdi [11].
@@
Let us elaborate on the main contributions and methodological advances of this article. To the best of our knowledge, this work appears to be the first to establish the existence of strong solutions for a nonlinearly coupled multilayered FSI problem, where the fluid is described by the Navier-Stokes equations. Moreover, it seems to be the first paper to address strong time-periodic solutions to a multilayered fluid-structure interaction problem. We believe that by a similar methodology as in the present paper, the corresponding initial value problem could be tackled to establish its local strong well-posedness and global strong well-posedness for small data. In fact, one would employ maximal $\mathrm{L}^{p}$-regularity instead of maximal periodic $\mathrm{L}^{p}$-regularity as in this manuscript. This is left to future study. The presence of the thick structural layer introduces substantial analytical challenges, requiring significant extensions beyond existing FSI theory. A key technical contribution is the bounded $\mathcal{H}^{\infty}$-calculus of the operator matrix associated with the damped wave system (see Subsection 5.2), which, to the best of our knowledge, is new and may be of independent interest. The presence of the thick layer also necessitates a further step in the lifting procedure, and the resulting multilayered fluid-structure operator is considerably more complex than in the case of a single thin layer as in [48]. In particular, to tackle the strongly coupled nature of the present multilayered problem, a more intricate decoupling approach tailored to the functional analytic setting of the damped wave problem is required. As far as we are aware, this is the first time such a technique has been employed in the context of boundary-coupled FSI problems with elastic structures, though it may prove useful in a broader class of FSI models. Furthermore, the damping in the thick structural layer renders the spectral analysis substantially more delicate. In particular, the determination of the essential spectrum requires a careful decomposition of the decoupled operator matrix.
@@
This article has the following outline. In Section 2, we settle some notation and make precise the function spaces used throughout this paper. In Section 3, we introduce the time-periodic multilayered FSI problem and present the main result on the existence of a strong time-periodic solution for small periodic forces. Section 4 is dedicated to the transformation to a fixed fluid domain. The central Section 5 discusses the linear theory for the time-periodic problem, including the damped wave problem, the reformulation in operator form, and the maximal $\mathrm{L}^{p}$-regularity and invertibility of the resulting operator matrix. In Section 6, we establish the nonlinear estimates and complete the proof of the main result.
2. Notation and function spaces
In this section, we make precise some pieces of notation and function spaces. In fact, the subscript f will be used to denote objects related to the fluids, while the subscript s indicates objects associated with the thick structural layer. On the other hand, the superscript F will be employed for variables on the moving fluid domains. Moreover, $\Omega$ will be used for three-dimensional domains, whereas $\Gamma$ will represent boundaries or interfaces. Concerning differential operators, we use $\Delta_{s}$ and $\nabla_{s}$ to designate the horizontal Laplacian and gradient, respectively, i.e.,
$\Delta_{s}=\partial_{1}^{2}+\partial_{2}^{2}\kern 5.0pt\text{and}\kern 5.0pt\nabla_{s}=\binom{\partial_{1}}{\partial_{2}}$.
For Banach spaces $\mathrm{X}$ and $\mathrm{Y}$, we denote by $\mathcal{L}(\mathrm{X},\mathrm{Y})$ the set of bounded linear operators from $\mathrm{X}$ to $\mathrm{Y}$, with the convention $\mathcal{L}(\mathrm{X})\coloneqq\mathcal{L}(\mathrm{X},\mathrm{X})$.
Next, let us introduce the function spaces used throughout the paper. For a sufficiently regular domain $G\subset\mathbb{R}^{n}$, $n\in\mathbb{N}$, as well as $m\in\mathbb{N}$ and $p\in[1,\infty]$, we use $\mathrm{L}^{p}(G)$ and $\mathrm{W}^{m,p}(G)$ to denote the Lebesgue and Sobolev spaces. We will use the subscript m to denote functions with average zero in $\mathrm{L}^{p}(G)$, i.e.,
$\mathrm{L}_{\mathrm{m}}^{p}(G)\coloneqq\left\{f\in\mathrm{L}^{p}(G):\int_{G}f\,\mathrm{d}x=0\right\}.$
Besides, $\mathrm{W}_{0}^{1,p}(G)$ denotes the functions in $\mathrm{W}^{1,p}(G)$ with zero trace on $\partial G$. Moreover, for $s\geq 0$, $p\in(1,\infty)$ and $q\in[1,\infty]$, we denote by $\mathrm{H}^{s,p}(G)$, $\mathrm{W}^{s,p}(G)$ and $\mathrm{B}_{pq}^{s}(G)$ the Bessel potential spaces, fractional Sobolev or Sobolev-Slobodeckij spaces and Besov spaces, respectively. Note that the Bessel potential spaces and Besov spaces $\mathrm{H}^{s,p}(\mathbb{R}^{n})$ and $\mathrm{B}_{pq}^{s}(\mathbb{R}^{n})$ on the whole space can be defined in terms of the Bessel potential and dyadic partitions, respectively, while we set
$\mathrm{W}^{s,p}(\mathbb{R}^{n})\coloneqq\left\{\begin{aligned} &\mathrm{H}^{s,p}(\mathbb{R}^{n})&&\kern 5.0pt\text{if}\kern 5.0pts\in\mathbb{N}_{0},\\ &\mathrm{B}_{pp}^{s}(\mathbb{R}^{n})&&\kern 5.0pt\text{if}\kern 5.0pts\notin\mathbb{N}_{0}.\end{aligned}\right$.
The respective spaces on $G$ can then be obtained via restriction. We do not provide more details here, but we refer to [65] instead for further background. Let us only observe that for $m\in\mathbb{N}$, $p\in(1,\infty)$, $q\in[1,\infty]$ and $\theta\in(0,1)$, we have the interpolation relations
$\mathrm{L}^{p}(G),\mathrm{W}^{m,p}(G)]_{\theta}=\mathrm{H}^{\theta m,p}(G)\kern 5.0pt\text{and}\kern 5.0pt(\mathrm{L}^{p}(G),\mathrm{W}^{m,p}(G))_{\theta,q}=\mathrm{B}_{pq}^{\theta m}(G),$
where $\cdot,\cdot]_{\theta}$ and $(\cdot,\cdot)_{\theta,q}$ represent the complex and real interpolation functor, respectively. For $s\geq 0$ and $p\in(1,\infty)$, we will also use the notation $\mathrm{W}_{\mathrm{m}}^{s,p}(G)\coloneqq\mathrm{W}^{s,p}(G)\cap\mathrm{L}_{\mathrm{m}}^{p}(G)$. In the context of continuous functions, we use the subscript c to indicate a compact support, while by $\mathrm{BUC}(G)$, we denote the bounded and uniformly continuous functions on $G$.
We conclude this section by introducing function spaces on moving domains, which will be required to rigorously define functions on the time-dependent fluid domain. To this end, we consider the reference fluid domain $\Omega_{\mathrm{f}}$ and the moving fluid domain $\Omega_{\mathrm{f}}(t)$, which will be introduced in Section 3. Moreover, we consider a $\mathrm{C}^{1}$-diffeomorphism $X$ mapping $\Omega_{\mathrm{f}}$ onto $\Omega_{\mathrm{f}}(t)$. This mapping will be constructed in Section 4 and is assumed to satisfy $X\in\mathrm{L}^{p}\bigl(0,T;\mathrm{W}^{2,q}(\Omega_{\mathrm{f}})^{3}\bigr)\cap\mathrm{W}^{2,p}\bigl(0,T;\mathrm{L}^{q}(\Omega_{\mathrm{f}})^{3}\bigr)$. For $p,q\in(1,\infty)$, we then define the following function spaces